\begin{thebibliography}{47}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Agrawal et~al.(2018)Agrawal, Turchi, and
  Negri}]{agrawal2018contextual}
Ruchit~Rajeshkumar Agrawal, Marco Turchi, and Matteo Negri. 2018.
\newblock Contextual handling in neural machine translation: Look behind, ahead
  and on both sides.
\newblock In \emph{21st Annual Conference of the European Association for
  Machine Translation}, pages 11--20.

\bibitem[{Bahdanau et~al.(2014)Bahdanau, Cho, and Bengio}]{bahdanau2014neural}
Dzmitry Bahdanau, Kyunghyun Cho, and Yoshua Bengio. 2014.
\newblock Neural machine translation by jointly learning to align and
  translate.
\newblock \emph{arXiv preprint arXiv:1409.0473}.

\bibitem[{Bao et~al.(2021)Bao, Zhang, Teng, Chen, and Luo}]{bao2021g}
Guangsheng Bao, Yue Zhang, Zhiyang Teng, Boxing Chen, and Weihua Luo. 2021.
\newblock G-transformer for document-level machine translation.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 3442--3455.

\bibitem[{Bawden et~al.(2018)Bawden, Sennrich, Birch, and
  Haddow}]{bawden2017evaluating}
Rachel Bawden, Rico Sennrich, Alexandra Birch, and Barry Haddow. 2018.
\newblock Evaluating discourse phenomena in neural machine translation.
\newblock In \emph{Proceedings of the 2018 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, pages 1304--1313.

\bibitem[{Cettolo et~al.(2017)Cettolo, Federico, Bentivogli, Niehues,
  St{\"u}ker, Sudoh, Yoshino, and Federmann}]{cettolo2017overview}
Mauro Cettolo, Marcello Federico, Luisa Bentivogli, Jan Niehues, Sebastian
  St{\"u}ker, Katsuitho Sudoh, Koichiro Yoshino, and Christian Federmann. 2017.
\newblock Overview of the iwslt 2017 evaluation campaign.
\newblock In \emph{Proceedings of the 14th International Workshop on Spoken
  Language Translation}, pages 2--14.

\bibitem[{Currey et~al.(2022)Currey, N{\u{a}}dejde, Pappagari, Mayer, Lauly,
  Niu, Hsu, and Dinu}]{currey2022mt}
Anna Currey, Maria N{\u{a}}dejde, Raghavendra Pappagari, Mia Mayer, Stanislas
  Lauly, Xing Niu, Benjamin Hsu, and Georgiana Dinu. 2022.
\newblock Mt-geneval: A counterfactual and contextual dataset for evaluating
  gender accuracy in machine translation.
\newblock \emph{arXiv preprint arXiv:2211.01355}.

\bibitem[{Farhad et~al.(2021)Farhad, Arkady, Magdalena, Ond{\v{r}}ej, Rajen,
  Vishrav, Costa-jussa, Cristina, Angela, Christian
  et~al.}]{farhad2021findings}
Akhbardeh Farhad, Arkhangorodsky Arkady, Biesialska Magdalena, Bojar
  Ond{\v{r}}ej, Chatterjee Rajen, Chaudhary Vishrav, Marta~R Costa-jussa,
  Espa{\~n}a-Bonet Cristina, Fan Angela, Federmann Christian, et~al. 2021.
\newblock Findings of the 2021 conference on machine translation (wmt21).
\newblock In \emph{Proceedings of the Sixth Conference on Machine Translation},
  pages 1--88. Association for Computational Linguistics.

\bibitem[{Fernandes et~al.(2021)Fernandes, Yin, Neubig, and
  Martins}]{fernandes2021measuring}
Patrick Fernandes, Kayo Yin, Graham Neubig, and Andr{\'e}~FT Martins. 2021.
\newblock Measuring and increasing context usage in context-aware machine
  translation.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 6467--6478.

\bibitem[{Garcia et~al.(2019)Garcia, Creus, and
  Espa{\~n}a-Bonet}]{garcia2019context}
Eva~Mart{\'\i}nez Garcia, Carles Creus, and Cristina Espa{\~n}a-Bonet. 2019.
\newblock Context-aware neural machine translation decoding.
\newblock In \emph{Proceedings of the Fourth Workshop on Discourse in Machine
  Translation (DiscoMT 2019)}, pages 13--23.

\bibitem[{Huo et~al.(2020)Huo, Herold, Gao, Dahlmann, Khadivi, and
  Ney}]{huo2020diving}
Jingjing Huo, Christian Herold, Yingbo Gao, Leonard Dahlmann, Shahram Khadivi,
  and Hermann Ney. 2020.
\newblock Diving deep into context-aware neural machine translation.
\newblock In \emph{Proceedings of the Fifth Conference on Machine Translation},
  pages 604--616.

\bibitem[{Jean et~al.(2017)Jean, Lauly, Firat, and Cho}]{jean2017does}
Sebastien Jean, Stanislas Lauly, Orhan Firat, and Kyunghyun Cho. 2017.
\newblock Does neural machine translation benefit from larger context?
\newblock \emph{arXiv preprint arXiv:1704.05135}.

\bibitem[{Jiang et~al.(2022)Jiang, Liu, Ma, Zhang, Yang, Huang, Sennrich,
  Cotterell, Sachan, and Zhou}]{jiang-etal-2022-blonde}
Yuchen Jiang, Tianyu Liu, Shuming Ma, Dongdong Zhang, Jian Yang, Haoyang Huang,
  Rico Sennrich, Ryan Cotterell, Mrinmaya Sachan, and Ming Zhou. 2022.
\newblock \href {https://doi.org/10.18653/v1/2022.naacl-main.111} {{BlonDe}: An
  automatic evaluation metric for document-level machine translation}.
\newblock In \emph{Proceedings of the 2022 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 1550--1565, Seattle, United States. Association for
  Computational Linguistics.

\bibitem[{Junczys-Dowmunt(2019)}]{junczys-dowmunt-2019-microsoft}
Marcin Junczys-Dowmunt. 2019.
\newblock \href {https://doi.org/10.18653/v1/W19-5321} {{M}icrosoft translator
  at {WMT} 2019: Towards large-scale document-level neural machine
  translation}.
\newblock In \emph{Proceedings of the Fourth Conference on Machine Translation
  (Volume 2: Shared Task Papers, Day 1)}, pages 225--233, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Jwalapuram et~al.(2019)Jwalapuram, Joty, Temnikova, and
  Nakov}]{jwalapuram2019evaluating}
Prathyusha Jwalapuram, Shafiq Joty, Irina Temnikova, and Preslav Nakov. 2019.
\newblock Evaluating pronominal anaphora in machine translation: An evaluation
  measure and a test suite.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pages 2964--2975.

\bibitem[{Kim et~al.(2019)Kim, Tran, and Ney}]{kim2019and}
Yunsu Kim, Duc~Thanh Tran, and Hermann Ney. 2019.
\newblock When and why is document-level context useful in neural machine
  translation?
\newblock In \emph{Proceedings of the Fourth Workshop on Discourse in Machine
  Translation (DiscoMT 2019)}, pages 24--34.

\bibitem[{Kudo(2018)}]{DBLP:conf/acl/Kudo18}
Taku Kudo. 2018.
\newblock \href {https://doi.org/10.18653/v1/P18-1007} {Subword regularization:
  Improving neural network translation models with multiple subword
  candidates}.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics, {ACL} 2018, Melbourne, Australia, July 15-20,
  2018, Volume 1: Long Papers}, pages 66--75. Association for Computational
  Linguistics.

\bibitem[{L{\"a}ubli et~al.(2018)L{\"a}ubli, Sennrich, and
  Volk}]{laubli2018has}
Samuel L{\"a}ubli, Rico Sennrich, and Martin Volk. 2018.
\newblock Has machine translation achieved human parity? a case for
  document-level evaluation.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 4791--4796.

\bibitem[{Li et~al.(2020)Li, Liu, Wang, Jiang, Xiao, Zhu, Liu, and
  Li}]{li2020does}
Bei Li, Hui Liu, Ziyang Wang, Yufan Jiang, Tong Xiao, Jingbo Zhu, Tongran Liu,
  and Changliang Li. 2020.
\newblock Does multi-encoder help? a case study on context-aware neural machine
  translation.
\newblock \emph{arXiv preprint arXiv:2005.03393}.

\bibitem[{Lison et~al.(2018)Lison, Tiedemann, and
  Kouylekov}]{lison2018opensubtitles2018}
Pierre Lison, J{\"o}rg Tiedemann, and Milen Kouylekov. 2018.
\newblock Opensubtitles2018: Statistical rescoring of sentence alignments in
  large, noisy parallel corpora.
\newblock In \emph{Proceedings of the Eleventh International Conference on
  Language Resources and Evaluation (LREC 2018)}.

\bibitem[{Liu et~al.(2020)Liu, Gu, Goyal, Li, Edunov, Ghazvininejad, Lewis, and
  Zettlemoyer}]{10.1162/tacl_a_00343}
Yinhan Liu, Jiatao Gu, Naman Goyal, Xian Li, Sergey Edunov, Marjan
  Ghazvininejad, Mike Lewis, and Luke Zettlemoyer. 2020.
\newblock \href {https://doi.org/10.1162/tacl_a_00343} {{Multilingual Denoising
  Pre-training for Neural Machine Translation}}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  8:726--742.

\bibitem[{Ma et~al.(2020)Ma, Zhang, and Zhou}]{ma2020simple}
Shuming Ma, Dongdong Zhang, and Ming Zhou. 2020.
\newblock A simple and effective unified encoder for document-level machine
  translation.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 3505--3511.

\bibitem[{Majumde et~al.(2022)Majumde, Lauly, Nadejde, Federico, and
  Dinu}]{majumde2022baseline}
Suvodeep Majumde, Stanislas Lauly, Maria Nadejde, Marcello Federico, and
  Georgiana Dinu. 2022.
\newblock A baseline revisited: Pushing the limits of multi-segment models for
  context-aware translation.
\newblock \emph{arXiv preprint arXiv:2210.10906}.

\bibitem[{Maruf and Haffari(2018)}]{maruf2018document}
Sameen Maruf and Gholamreza Haffari. 2018.
\newblock Document context neural machine translation with memory networks.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1275--1284.

\bibitem[{Maruf et~al.(2019)Maruf, Martins, and Haffari}]{maruf2019selective}
Sameen Maruf, Andr{\'e}~FT Martins, and Gholamreza Haffari. 2019.
\newblock Selective attention for context-aware neural machine translation.
\newblock In \emph{Proceedings of the 2019 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 3092--3102.

\bibitem[{Miculicich et~al.(2018)Miculicich, Ram, Pappas, and
  Henderson}]{miculicich2018document}
Lesly Miculicich, Dhananjay Ram, Nikolaos Pappas, and James Henderson. 2018.
\newblock Document-level neural machine translation with hierarchical attention
  networks.
\newblock \emph{arXiv preprint arXiv:1809.01576}.

\bibitem[{M{\"u}ller et~al.(2018)M{\"u}ller, Gonzales, Voita, and
  Sennrich}]{muller2018large}
Mathias M{\"u}ller, Annette~Rios Gonzales, Elena Voita, and Rico Sennrich.
  2018.
\newblock A large-scale test set for the evaluation of context-aware pronoun
  translation in neural machine translation.
\newblock In \emph{Proceedings of the Third Conference on Machine Translation:
  Research Papers}, pages 61--72.

\bibitem[{Nguyen et~al.(2021)Nguyen, Murray, and Chiang}]{nguyen2021data}
Toan~Q Nguyen, Kenton Murray, and David Chiang. 2021.
\newblock Data augmentation by concatenation for low-resource translation: A
  mystery and a solution.
\newblock \emph{IWSLT 2021}, page 287.

\bibitem[{Ott et~al.(2019)Ott, Edunov, Baevski, Fan, Gross, Ng, Grangier, and
  Auli}]{ott2019fairseq}
Myle Ott, Sergey Edunov, Alexei Baevski, Angela Fan, Sam Gross, Nathan Ng,
  David Grangier, and Michael Auli. 2019.
\newblock fairseq: A fast, extensible toolkit for sequence modeling.
\newblock In \emph{Proceedings of the 2019 Conference of the North American
  Chapter of the Association for Computational Linguistics (Demonstrations)},
  pages 48--53.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and
  Zhu}]{papineni2002bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock Bleu: a method for automatic evaluation of machine translation.
\newblock In \emph{Proceedings of the 40th annual meeting of the Association
  for Computational Linguistics}, pages 311--318.

\bibitem[{Petrick et~al.(2022)Petrick, Rosendahl, Herold, and
  Ney}]{petrick2022locality}
Frithjof Petrick, Jan Rosendahl, Christian Herold, and Hermann Ney. 2022.
\newblock Locality-sensitive hashing for long context neural machine
  translation.
\newblock In \emph{Proceedings of the 19th International Conference on Spoken
  Language Translation (IWSLT 2022)}, pages 32--42.

\bibitem[{Post(2018)}]{post2018call}
Matt Post. 2018.
\newblock A call for clarity in reporting bleu scores.
\newblock In \emph{Proceedings of the Third Conference on Machine Translation:
  Research Papers}, pages 186--191.

\bibitem[{Sennrich et~al.(2016)Sennrich, Haddow, and
  Birch}]{sennrich2016neural}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016.
\newblock Neural machine translation of rare words with subword units.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1715--1725.

\bibitem[{Snover et~al.(2006)Snover, Dorr, Schwartz, Micciulla, and
  Makhoul}]{snover2006study}
Matthew Snover, Bonnie Dorr, Richard Schwartz, Linnea Micciulla, and John
  Makhoul. 2006.
\newblock A study of translation edit rate with targeted human annotation.
\newblock In \emph{Proceedings of the 7th Conference of the Association for
  Machine Translation in the Americas: Technical Papers}, pages 223--231.

\bibitem[{Sun et~al.(2022)Sun, Wang, Zhou, Zhao, Huang, Chen, and
  Li}]{sun2022rethinking}
Zewei Sun, Mingxuan Wang, Hao Zhou, Chengqi Zhao, Shujian Huang, Jiajun Chen,
  and Lei Li. 2022.
\newblock Rethinking document-level neural machine translation.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL 2022}, pages 3537--3548.

\bibitem[{Thai et~al.(2022)Thai, Karpinska, Krishna, Ray, Inghilleri, Wieting,
  and Iyyer}]{thai2022exploring}
Katherine Thai, Marzena Karpinska, Kalpesh Krishna, Bill Ray, Moira Inghilleri,
  John Wieting, and Mohit Iyyer. 2022.
\newblock Exploring document-level literary machine translation with parallel
  paragraphs from world literature.
\newblock \emph{arXiv preprint arXiv:2210.14250}.

\bibitem[{Tiedemann and Scherrer(2017)}]{tiedemann2017neural}
J{\"o}rg Tiedemann and Yves Scherrer. 2017.
\newblock Neural machine translation with extended context.
\newblock In \emph{Proceedings of the Third Workshop on Discourse in Machine
  Translation}, pages 82--92.

\bibitem[{Tu et~al.(2018)Tu, Liu, Shi, and Zhang}]{tu2018learning}
Zhaopeng Tu, Yang Liu, Shuming Shi, and Tong Zhang. 2018.
\newblock Learning to remember translation history with a continuous cache.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  6:407--420.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, \L~ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock \href
  {https://proceedings.neurips.cc/paper/2017/file/3f5ee243547dee91fbd053c1c4a845aa-Paper.pdf}
  {Attention is all you need}.
\newblock In \emph{Advances in Neural Information Processing Systems},
  volume~30. Curran Associates, Inc.

\bibitem[{Voita et~al.(2019)Voita, Sennrich, and Titov}]{voita2019good}
Elena Voita, Rico Sennrich, and Ivan Titov. 2019.
\newblock \href {http://arxiv.org/abs/1905.05979} {When a good translation is
  wrong in context: Context-aware machine translation improves on deixis,
  ellipsis, and lexical cohesion}.

\bibitem[{Voita et~al.(2018)Voita, Serdyukov, Sennrich, and
  Titov}]{voita2018context}
Elena Voita, Pavel Serdyukov, Rico Sennrich, and Ivan Titov. 2018.
\newblock Context-aware neural machine translation learns anaphora resolution.
\newblock \emph{arXiv preprint arXiv:1805.10163}.

\bibitem[{Wang et~al.(2017)Wang, Tu, Way, and Liu}]{wang2017exploiting}
Longyue Wang, Zhaopeng Tu, Andy Way, and Qun Liu. 2017.
\newblock Exploiting cross-sentence context for neural machine translation.
\newblock \emph{arXiv preprint arXiv:1704.04347}.

\bibitem[{Wong et~al.(2020)Wong, Maruf, and Haffari}]{wong2020contextual}
KayYen Wong, Sameen Maruf, and Gholamreza Haffari. 2020.
\newblock Contextual neural machine translation improves translation of
  cataphoric pronouns.
\newblock \emph{arXiv preprint arXiv:2004.09894}.

\bibitem[{Xiong et~al.(2019)Xiong, He, Wu, and Wang}]{xiong2019modeling}
Hao Xiong, Zhongjun He, Hua Wu, and Haifeng Wang. 2019.
\newblock Modeling coherence for discourse neural machine translation.
\newblock In \emph{Proceedings of the AAAI conference on artificial
  intelligence}, volume~33, pages 7338--7345.

\bibitem[{Yang et~al.(2022)Yang, Yin, Yang, Ma, Huang, Zhang, Wei, and
  Li}]{yang2022gtrans}
Jian Yang, Yuwei Yin, Liqun Yang, Shuming Ma, Haoyang Huang, Dongdong Zhang,
  Furu Wei, and Zhoujun Li. 2022.
\newblock Gtrans: Grouping and fusing transformer layers for neural machine
  translation.
\newblock \emph{IEEE/ACM Transactions on Audio, Speech, and Language
  Processing}.

\bibitem[{Zhang et~al.(2018)Zhang, Luan, Sun, Zhai, Xu, Zhang, and
  Liu}]{zhang2018improving}
Jiacheng Zhang, Huanbo Luan, Maosong Sun, Feifei Zhai, Jingfang Xu, Min Zhang,
  and Yang Liu. 2018.
\newblock Improving the transformer translation model with document-level
  context.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 533--542.

\bibitem[{Zhang et~al.(2020)Zhang, Chen, Ge, and Fan}]{zhang-etal-2020-long}
Pei Zhang, Boxing Chen, Niyu Ge, and Kai Fan. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.81} {Long-short
  term masking transformer: A simple but effective baseline for document-level
  neural machine translation}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 1081--1087, Online. Association
  for Computational Linguistics.

\bibitem[{Zheng et~al.(2021)Zheng, Yue, Huang, Chen, and
  Birch}]{zheng2021towards}
Zaixiang Zheng, Xiang Yue, Shujian Huang, Jiajun Chen, and Alexandra Birch.
  2021.
\newblock Towards making the most of context in neural machine translation.
\newblock In \emph{Proceedings of the Twenty-Ninth International Conference on
  International Joint Conferences on Artificial Intelligence}, pages
  3983--3989.

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt]{article}

\usepackage{lipsum}  
\usepackage{booktabs}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsbsy}
\usepackage{xspace}
\usepackage{multirow}
\DeclareMathOperator*{\argmax}{argmax}
\usepackage{url}

\newcommand\BLEU{\textsc{Bleu}\xspace}
\newcommand\TER{\textsc{Ter}\xspace}

\usepackage[]{ACL2023}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\title{On Search Strategies for Document-Level Neural Machine Translation}

\author{
Christian Herold \qquad Hermann Ney \\
Human Language Technology and Pattern Recognition Group \\
Computer Science Department\\
RWTH Aachen University \\
D-52056 Aachen, Germany \\
{\tt \{herold|ney\}@cs.rwth-aachen.de}
}

\begin{document}
\maketitle
\begin{abstract}

Compared to sentence-level systems, document-level neural machine translation (NMT) models produce a more consistent output across a document and are able to better resolve ambiguities within the input.
There are many works on document-level NMT, mostly focusing on modifying the model architecture or training strategy to better accommodate the additional context-input.
On the other hand, in most works, the question on how to perform search with the trained model is scarcely discussed, sometimes not mentioned at all.
In this work, we aim to answer the question how to best utilize a context-aware translation model in decoding.
We start with the most popular document-level NMT approach and compare different decoding schemes, some from the literature and others proposed by us.
In the comparison, we are using both, standard automatic metrics, as well as specific linguistic phenomena on three standard document-level translation benchmarks.
We find that most commonly used decoding strategies perform similar to each other and that higher quality context information has the potential to further improve the translation.
\end{abstract}

\section{Introduction}

Neural machine translation (NMT) \cite{bahdanau2014neural, vaswani2017attention} is widely adopted and produces excellent translations for many domains and language pairs.
However, when these automatic translations are evaluated on the document level, they reveal shortcomings when it comes to consistency in style, entity-translation or correct inference of the gender, among other things \cite{laubli2018has, muller2018large, thai2022exploring}.
Document-level NMT aims to resolve these shortcomings by taking the context of a sentence into account during translation.
There exist many works on the topic of document-level NMT, proposing various changes to the standard transformer \cite{vaswani2017attention} architecture and training criteria to improve context incorporation and consequently translation quality.
However, while the modeling and training aspects are covered in great detail in these works, the exact decoding strategy is often not very clearly described and sometimes not mentioned at all.

In this work, we head out to answer the question, which decoding strategy is most beneficial for document-level NMT systems.
We compare all commonly used strategies, as well as some additional ones, on three standard document-level translation benchmarks.
We find that most of the analyzed decoding strategies perform similar to each other.
Also, higher quality context information can lead to better translations in certain scenarios.

\section{Related Work}

The earliest approaches to document-level NMT simply concatenate consecutive sentences without any further changes to the architecture compared to the sentence-level systems \cite{tiedemann2017neural, agrawal2018contextual}.
Later, some changes were made to the vanilla transformer architecture, like segment embeddings \cite{ma2020simple} or attention masking \cite{zhang-etal-2020-long, petrick2022locality} and a move was made towards translating longer segments \cite{junczys-dowmunt-2019-microsoft, 10.1162/tacl_a_00343, zheng2021towards, bao2021g, sun2022rethinking}.
Other works employ a separate encoder to include the additional context on the source side \cite{jean2017does, bawden2017evaluating, zhang2018improving, voita2018context} or make use of the context in a post-editing fashion \cite{voita2019good, xiong2019modeling}.
Further approaches include the usage of a cache \cite{wang2017exploiting, maruf2018document, tu2018learning} or hierarchical attention networks \cite{miculicich2018document, maruf2019selective, wong2020contextual}.
Recently, several works have concluded that the simple concatenation approach used with the vanilla transformer architecture performs as good - if not better - than more complicated approaches that modify the model structure \cite{sun2022rethinking, majumde2022baseline}.
Since we also observed this in our internal comparisons, we decided to focus on this simple approach for our analysis in this work. 

Several works made the argument that the improvements seen in automatic metric scores for document-level NMT systems are from regularization effects rather than from utilizing the additional context information \cite{kim2019and, li2020does, nguyen2021data}.
In order to better asses the improvements gained by document-level NMT, several targeted test suites have been released \cite{muller2018large, bawden2017evaluating, voita2019good, jwalapuram2019evaluating}.
However, all of these are based on just scoring contrastive examples without actually translating anything.
Recently, \citet{jiang-etal-2022-blonde} and \citet{currey2022mt} have released frameworks that allow to score MT systems on their ability to generate contextually correct translations.\footnote{The framework by \citet{currey2022mt} was not yet made available when we conducted our experiments.}

\section{Search Strategies}
\label{sec:methodology}

Training a document-level NMT system that takes the last $k$ sentences as context is straightforward using the standard concatenation strategy \cite{tiedemann2017neural}.
Given some document level training data $(F_n, E_n), n=1, ..., N$, where $(F_n, E_n)$ denotes the $n$-th source-target sentence pair, during training we optimize the parameters $\Theta$ of the model towards
\begin{equation*}
    \hat{\Theta} = \argmax_{\Theta}\left\{ \sum_n \log p_{\Theta}(E_{n-k}^{n}| F_{n-k}^{n}) \right\}.
\end{equation*}
Here, $E_{n-k}^{n}$ denotes the concatenation of the sentences $E_{n-k}, ..., E_n$.

During search, given a document $F_1^M$, we want to find the best translation $\hat{E}_1^M$ according to the model.
Of course, exact search can not be performed and different works have used different methods to generate a translation:
\begin{description}
    \item[\textit{full segment}]\cite{10.1162/tacl_a_00343, bao2021g, sun2022rethinking}: we split the document into non-overlapping parts $F_1^{k}, F_{k+1}^{2k}, ..., F_{M-k}^{M}$ and translate each part separately using
    \begin{equation}
        \hat{E}_{i-k}^{i} = \argmax_{E_{i-k}^{i}}\left\{p(E_{i-k}^{i}|F_{i-k}^{i}) \right\},
        \label{eq:1}
    \end{equation}
    which is approximated using standard beam search on the token level.
    \item[\textit{last sentence}]\cite{bawden2017evaluating, agrawal2018contextual, zhang-etal-2020-long, petrick2022locality, majumde2022baseline}: we split the document into overlapping parts $..., F_{i-k}^{i}, F_{i-k+1}^{i+1}, ...$ and translate each part separately using Equation \ref{eq:1}. From each translated part we choose only the last sentence to get one translation for every sentence in the document.
    \item[\textit{first sentence}]\cite{zhang-etal-2020-long}: similar to \textit{last sentence}, but from each translated part we choose only the first sentence to get one translation for every sentence in the document.
    \item[\textit{2-pass decoding}]\cite{maruf2018document, maruf2019selective, voita2019good, xiong2019modeling}: we first generate a translation $\tilde{E}_1^M$ of the document using a sentence-level NMT system. Then, the final hypothesis $\hat{E}_i$ for each sentence $F_i$ is created using 
    \begin{equation*}
        \hat{E}_{i} = \argmax_{E_{i}}\left\{p(E_{i}|F_{i-k}^{i}, \tilde{E}_{i-k}^{i-1}) \right\}.
    \end{equation*}
    \item[\textit{doc-trans}]\cite{miculicich2018document, voita2019good, garcia2019context, fernandes2021measuring}: we generate the translation sentence by sentence, meaning
    \begin{equation*}
        \hat{E}_{1} = \argmax_{E_{1}}\left\{p(E_{1}|F_{1}) \right\},
    \end{equation*}
    \begin{equation*}
        \hat{E}_{2} = \argmax_{E_{2}}\left\{p(E_{2}|F_{1}^{2}, \hat{E_1}) \right\},
    \end{equation*}
        \begin{equation*}
        ...
    \end{equation*}
    \item[\textit{doc-trans (beam)}]: similar to \textit{doc-trans}, but instead of keeping just the best context $\hat{E}_{1}^{i-1}$, we keep the top-$h$ candidates and prune them after each step $i$, analogous to beam search on the token level. $h=12$ for all our experiments, the same as our token-level beam-size.
    \item[\textit{cheating}]: this is just used as a tool for analysis. The translation of each sentence $F_i$ is created using the true target reference $\mathring{E}_{1}^{M}$ as context
    \begin{equation*}
        \hat{E}_{i} = \argmax_{E_{i}}\left\{p(E_{i}|F_{i-k}^{i}, \mathring{E}_{i-k}^{i-1}) \right\}.
    \end{equation*}
        \item[\textit{no context}]: this is just used as a tool for analysis. The translation of each sentence $F_i$ is created using no context information at all
    \begin{equation*}
        \hat{E}_{i} = \argmax_{E_{i}}\left\{p(E_{i}|F_{i}) \right\}.
    \end{equation*}
\end{description}

The different search strategies also have a different computational cost associated with them.
The biggest factor regarding the decoding cost is the number of forward passes through the model, specifically the decoder, that we have to do.
We list the computational costs for the different decoding approaches in Table \ref{tab:decoding_cost} under the assumption that the document consists of $N$ sentences with average sentence length $L$ and the model uses $k-1$ sentences as context.
\begin{table}[]
\centering
\begin{tabular}{l|r}
\toprule
 & \multicolumn{1}{c}{Cost} \\ \hline
\textbf{sentence-level} & $\mathcal{O}(NL)$ \\ \hline
\textbf{document-level} &  \\
\textit{full segment} & $\mathcal{O}(NL)$ \\
\textit{last sentence} & $\mathcal{O}(NLk)$ \\
\textit{first sentence} & $\mathcal{O}(NLk)$ \\
\textit{2-pass decoding} & $\mathcal{O}(2NL)$ \\
\textit{doc trans} & $\mathcal{O}(NL)$ \\
\textit{doc trans (beam)} & $\mathcal{O}(NLh)$ \\ \bottomrule
\end{tabular}
\caption{Computational cost of decoding (=number of forward passes through the decoder) for each of the search strategies described above. $h$ denotes the sentence-level beam size.}
\label{tab:decoding_cost}
\end{table}
Please note that the decoding time might follow a different dependence than the cost in the above table, since it heavily depends on the available hardware.
For example, \textit{doc trans} and \textit{doc trans (beam)} might have the same decoding time, if we have enough computational resources available, since the additional computations in \textit{doc trans (beam)} can all be done in parallel.

\section{Experiments}

We perform experiments on three document-level translation benchmarks, called \textbf{NEWS} (En$\to$De), \textbf{TED} (En$\to$It) and \textbf{OS} (En$\to$De).
For the details regarding data conditions and preparation, as well as model training, we refer to Appendix \ref{sec:appendix}.
For the context-aware systems, we concatenate 3 adjacent sentences (i.e. $k=3$) using a special token \texttt{<sep>}.
For the two En$\to$De tasks, we also evaluate the systems on the ContraPro test set \cite{muller2018large}.
Instead of scoring and ranking the contrastive examples in ContraPro, as the authors have originally envisioned, we translate the source side to calculate \BLEU and \TER as well as to score the pronoun translations according to Section \ref{subsec:eval}.
We can not evaluate the \textit{full segment} search strategy on ContraPro, because the sentences are not adjacent since they come from different documents.

\subsection{Evaluating Pronoun Translation}
\label{subsec:eval}

% table 1 moved for formatting reasons
\begin{table}[]
\centering
\begin{tabular}{l|r|r|r}
\toprule
 & \multicolumn{1}{c|}{NEWS} & \multicolumn{1}{c|}{TED} & \multicolumn{1}{c}{OS} \\ \hline
\textbf{sentence-level} &  &  &  \\
ref & 4.61 & 4.15 & 2.97 \\
hyp & 1.63 & 1.56 & 1.48 \\ \hline
\textbf{document-level} &  &  &  \\
ref & 4.46 & 3.96 & 2.70 \\
hyp \textit{no context} & 1.64 & 1.53 & 1.47 \\
hyp \textit{full segment} & 1.62 & 1.50 & 1.41 \\
hyp \textit{last sentence} & 1.61 & 1.48 & 1.42 \\
hyp \textit{first sentence} & 1.63 & 1.52 & 1.48 \\
hyp \textit{2-pass decoding} & 1.68 & 1.49 & 1.48 \\
hyp \textit{doc trans} & 1.67 & 1.48 & 1.42 \\
hyp \textit{doc trans (beam)} & 1.67 & 1.48 & 1.41 \\
hyp \textit{cheating} & 1.69 & 1.53 & 1.56 \\ \bottomrule
\end{tabular}
\caption{Perplexity values on the test set for different search strategies.}
\label{tab:ppls}
\end{table}

As further analysis, we measure how well ambiguous pronouns are handled when translating from English to German.
Regarding gender, the English third-person pronoun \lq{}it\rq{} (and its other forms), can be translated to the German words \lq{}er\rq{}, \lq{}sie\rq{} or \lq{}es\rq{}, depending on which noun it refers to.
On the other hand, ambiguities in the formality come from second-person pronouns.
For example, the English word \lq you\rq can be translated to \lq{}sie\rq{} or \lq{}du\rq{} depending if we are in a formal setting or not.
To report accuracies for pronoun (3 classes: male/female/neuter) and formality (2 classes: formal/informal) translation, we extend the BlonDe metric created by \citet{jiang-etal-2022-blonde}\footnote{Our extension can be found in this fork: \url{https://github.com/christian3141/BlonDe}}.
\begin{table*}[t]
\centering
\begin{tabular}{l|rrrr|rr|rrrr}
\toprule
 & \multicolumn{4}{c|}{NEWS} & \multicolumn{2}{c|}{TED} & \multicolumn{4}{c}{OS} \\ \cline{2-11} 
 & \multicolumn{2}{c|}{test} & \multicolumn{2}{c|}{ConPro} & \multicolumn{2}{c|}{test} & \multicolumn{2}{c|}{test} & \multicolumn{2}{c}{ConPro} \\
 & \multicolumn{1}{c}{\BLEU} & \multicolumn{1}{c|}{\TER} & \multicolumn{1}{c}{\BLEU} & \multicolumn{1}{c|}{\TER} & \multicolumn{1}{c}{\BLEU} & \multicolumn{1}{c|}{\TER} & \multicolumn{1}{c}{\BLEU} & \multicolumn{1}{c|}{\TER} & \multicolumn{1}{c}{\BLEU} & \multicolumn{1}{c}{\TER} \\ \hline
\textbf{sentence-level} &  & \multicolumn{1}{r|}{} &  &  &  &  &  & \multicolumn{1}{r|}{} &  &  \\
\textit{external} & $^{\dagger}$32.3 & \multicolumn{1}{r|}{-} & - & - & $^{\ddagger}$33.4 & - & *37.3 & \multicolumn{1}{r|}{-} & *30.5 & - \\
\textit{ours} & 32.8 & \multicolumn{1}{r|}{49.0} & 18.4 & 65.5 & 34.2 & 46.3 & 37.1 & \multicolumn{1}{r|}{43.8} & 29.7 & 52.8 \\ \hline
\textbf{document-level} &  & \multicolumn{1}{r|}{} &  &  &  &  &  & \multicolumn{1}{r|}{} &  &  \\
\textit{no context} & 33.4 & \multicolumn{1}{r|}{48.5} & 18.6 & 65.6 & 34.0 & 46.7 & 36.9 & \multicolumn{1}{r|}{44.5} & 29.4 & 53.2 \\
\textit{full segment} & 33.4 & \multicolumn{1}{r|}{48.6} & - & - & 34.3 & 46.3 & 38.2 & \multicolumn{1}{r|}{43.9} & - & - \\
\textit{last sentence} & 33.4 & \multicolumn{1}{r|}{48.3} & 19.7 & 63.4 & 34.7 & 45.9 & 37.8 & \multicolumn{1}{r|}{43.9} & 31.4 & 51.5 \\
\textit{first sentence} & 33.4 & \multicolumn{1}{r|}{48.6} & 18.8 & 65.6 & 34.1 & 46.3 & 37.8 & \multicolumn{1}{r|}{44.1} & 29.5 & 53.1 \\
\textit{2-pass decoding} & 32.8 & \multicolumn{1}{r|}{48.6} & 19.5 & 63.9 & 34.5 & 46.2 & 37.4 & \multicolumn{1}{r|}{44.4} & 31.1 & 51.8 \\
\textit{doc trans} & 33.0 & \multicolumn{1}{r|}{48.3} & 19.8 & 63.3 & 34.6 & 46.0 & 37.7 & \multicolumn{1}{r|}{44.0} & 31.4 & 51.3 \\
\textit{doc trans (beam)} & 33.0 & \multicolumn{1}{r|}{48.3} & 19.7 & 63.4 & 34.5 & 46.0 & 38.3 & \multicolumn{1}{r|}{43.8} & 31.3 & 51.5 \\
\textit{cheating} & 32.2 & \multicolumn{1}{r|}{49.4} & 19.3 & 65.1 & 34.1 & 46.6 & 39.6 & \multicolumn{1}{r|}{42.9} & 33.3 & 49.1 \\ \bottomrule
\end{tabular}
\caption{\BLEU and \TER scores (in percent) for the different tasks and decoding strategies. External baselines are from $^{\dagger}$ \citet{kim2019and}, $^{\ddagger}$ \citet{yang2022gtrans} and *\citet{huo2020diving}.}
\label{tab:bleu}
\end{table*}
\begin{table}[t]
\centering
\begin{tabular}{l|r|rr}
\toprule
 & \multicolumn{1}{c|}{NEWS} & \multicolumn{2}{c}{OS} \\ \cline{2-4} 
 & \multicolumn{1}{c|}{ConPro} & \multicolumn{1}{c|}{test} & \multicolumn{1}{c}{ConPro} \\
 & \multicolumn{1}{c|}{gender} & \multicolumn{1}{c|}{style} & \multicolumn{1}{c}{gender} \\ \hline
\textbf{sentence-level} & 45.3 & \multicolumn{1}{r|}{59.4} & 41.4 \\ \hline
\textbf{document-level} &  & \multicolumn{1}{r|}{} &  \\
\textit{no context} & 45.9 & \multicolumn{1}{r|}{59.7} & 42.3 \\
\textit{full segment} & - & \multicolumn{1}{r|}{60.8} & - \\
\textit{last sentence} & 56.1 & \multicolumn{1}{r|}{60.3} & 66.5 \\
\textit{first sentence} & 44.9 & \multicolumn{1}{r|}{59.2} & 43.0 \\
\textit{2-pass decoding} & 55.6 & \multicolumn{1}{r|}{59.9} & 65.1 \\
\textit{doc trans} & 56.1 & \multicolumn{1}{r|}{58.7} & 66.4 \\
\textit{doc trans (beam)} & 56.1 & \multicolumn{1}{r|}{60.6} & 66.3 \\
\textit{cheating} & 63.3 & \multicolumn{1}{r|}{73.2} & 73.7 \\ \bottomrule
\end{tabular}
\caption{F1 scores (in percent) for pronoun translation on different test sets.}
\label{tab:f1}
\end{table}
First, we expand the framework to work for German references, by including German NER and POS taggers\footnote{\url{https://spacy.io/models/de}} as well as including German pronoun mappings.
For the gender category, we mostly follow \citet{jiang-etal-2022-blonde}, but additionally require that a corresponding pronoun must also be present in the source sentence.\footnote{In ContraPro, we find 5011/4085/4817 examples for male/female/neuter respectively. The difference to the 4000/4000/4000 reported by \cite{muller2018large} means that in some cases we count multiple occurrences in a single example.}
For the style category, we take into account examples where a second person pronoun appears in the source sentence, and a corresponding formal or informal pronoun appears in the reference.\footnote{In the OS test set, we count 416 and 605 examples for formal and informal examples respectively.}

\subsection{Perplexities}

% table 2 and 3 moved for formatting reasons

First, we compare the perplexities of the hypotheses from the different search strategies, which are listed in Table \ref{tab:ppls}.
The first thing to note is, that the reference has a much higher perplexity than all hypotheses, which is commonly seen for NMT systems.
All document-level search strategies result in different hypotheses, which however have a similar perplexity score.
Surprisingly, the \textit{cheating} setting generates the worst translation perplexity-wise, even worse than using \textit{no context}.
This might be related to the observation, that the reference has a worse perplexity than any hypothesis, which is rather a modelling error than a search error.

\subsection{Automatic Metrics}

Next, we evaluate the hypotheses based on the common automatic metrics \BLEU and \TER.
The results are shown in Table \ref{tab:bleu}.
The hypotheses created with \textit{no context} seem to have the same quality as the sentence-level baseline.
Surprisingly, the true reference as context does not improve performance on the NEWS and TED test sets.
This indicates that the improvements seen on these test sets for the document-level system might not be related to better context incorporation.
On the contrary, the OS system creates the best hypothesis with the true reference as context.
All the actual decoding strategies give similar performance in terms of \BLEU and \TER with \textit{2-pass decoding} being a little bit behind.
A special case is the \textit{first sentence} strategy, which performs quite well on the standard test sets but poorly on ContraPro.
This is, because ContraPro is designed in a way that the left side context is more important for translation than the right side.

Finally, we analyze the quality of the pronoun translation as discussed in Section \ref{subsec:eval}.
In principle, we could calculate the F1 score for both, gender and formality, on all En$\to$De test sets.
However, we discard the cases where one or more classes have less than 100 examples.
This leaves us with the three test sets depicted in Table \ref{tab:f1}.
As a sanity check, we also report the ContraPro accuracies calculated from scoring the contrastive references as described in \cite{muller2018large}.
They are 48.2/45.8 for sentence-level and 68.2/82.2 for document-level for NEWS/OS respectively.
That means, with just scoring, we overestimate the capabilities of the system, but the trend is still consistent.\footnote{The precision and recall are roughly the same for the F1 scores reported in Table \ref{tab:f1}.}
Using the true reference leads to the best results in all cases.
\textit{no context} and \textit{first sentence} leaves us with sentence-level performance on the gender tasks, while all other decoding strategies perform similarly.
For the formality, none of the methods can significantly outperform the sentence level system, although the \textit{cheating} experiment shows that the system could do better if a better context information is provided.
This might be, because segments of 3 sentences are too short to reliably detect if a setting is formal or informal, without access to the true reference.

\section{Conclusion}

In this work, we analyze decoding strategies for document-level NMT systems.
Using the most popular document-level translation approach, we compare different search strategies found in the literature against methods developed by us.
We find that most of the commonly used decoding strategies result in similar performance, both in terms of common automatic metrics, as well as on specific pronoun evaluation tasks.
Therefore, we conclude that it is important to include the context information during decoding, but the exact way in which to do this is not as important.
Also, we find that the document-level systems could actually profit from higher quality context information, in situations where this context is most relevant for translation.

\section*{Acknowledgements}
This work was partially supported by the project HYKIST funded by the German Federal Ministry of Health on the basis of a decision of the German Federal Parliament (Bundestag) under funding ID ZMVI1-2520DAT04A, and by NeuroSys which, as part of the initiative “Clusters4Future”, is funded by the Federal Ministry of Education and Research BMBF (03ZU1106DA).

\section*{Limitations}
In this work, we limit our experiments to the most commonly used document-level system architecture and training criterion.
Other approaches exist, which might exhibit a different behavior in decoding.
Two out of the three document-level translation tasks we use in this work are low resource with less than 500k sentence-pairs as training data.
We chose these tasks due to computational limitations and to be better comparable to other works, but higher resource scenarios are more realistic for actual applications.
We limit the analysis of pronoun translation to the English-German language pair.
Also, there are other aspects of document-level NMT, like consistent translation of entities, which we did not consider in our analysis.

% \section*{Acknowledgements}

% Entries for the entire Anthology, followed by custom entries
\bibliography{anthology,custom}
\bibliographystyle{acl_natbib}
% Limitations
% \appendix

% \section{Example Appendix}
% \label{sec:appendix}

% This is a section in the appendix.

\clearpage

\appendix

\section{Appendix}
\label{sec:appendix}

For the \textbf{NEWS En$\to$De} task, the parallel training data (around 300k sentence pairs, news-domain) comes from the \texttt{NewsCommentaryV14} corpus\footnote{\url{https://data.statmt.org/news-commentary/v14/}}.
As validation/test set we use the WMT \texttt{newstest2015}/\texttt{newstest2018} test sets from the WMT news translation tasks \cite{farhad2021findings}.
For the \textbf{TED En$\to$It} task, the parallel training data (around 200k sentence pairs, scientific-talks-domain) comes from the IWSLT17 Multilingual Task \cite{cettolo2017overview}.
As validation set we use the concatenation of \texttt{IWSLT17.TED.dev2010} and \texttt{IWSLT17.TED.tst2010} and as test set we use \texttt{IWSLT17.TED.tst2017.mltlng}.
For the \textbf{OS En$\to$De} task, the parallel training data (around 22.5M sentence pairs, subtitle-domain) comes from the \texttt{OpenSubtitlesV2018} corpus \cite{lison2018opensubtitles2018}.
We use the same train/validation/test splits as \citet{huo2020diving} and additionally remove all segments that are used in the ContraPro test suite \cite{muller2018large} from the training data.
The data statistics for all tasks can be found in Table \ref{tab:data}.

\begin{table}[h!]
\centering
\begin{tabular}{l|l|r|r}
\toprule
\multicolumn{1}{c|}{task} & \multicolumn{1}{c|}{dataset} & \multicolumn{1}{c|}{\# sent.} & \multicolumn{1}{c}{\# doc.} \\ \hline
NEWS & train & 330k & 8.5k \\
 & valid & 2.2k & 81 \\
 & test & 3k & 122 \\
 & ContraPro & 12k & 12k \\ \hline
TED & train & 232k & 1.9k \\
 & valid & 2.5k & 19 \\
 & test & 1.1k & 10 \\ \hline
OS & train & 22.5M & 29.9k \\
 & valid & 3.5k & 5 \\
 & test & 3.8k & 5 \\
 & ContraPro & 12k & 12k \\ \bottomrule
\end{tabular}
\caption{Data statistics for the different document-level translation tasks.}
\label{tab:data}
\end{table}

Since in the original release of ContraPro only left side context is provided, we extract the right side context ourselves from \texttt{OpenSubtitlesV2018} based on the meta-information of the segments.

We tokenize the data using byte-pair-encoding \cite{sennrich2016neural, DBLP:conf/acl/Kudo18} with 15k joint merge operations (32k for OS En$\to$De).
The models are implemented using the fairseq toolkit \cite{ott2019fairseq} following the transformer base architecture \cite{vaswani2017attention} with dropout 0.3 and label-smoothing 0.2 for \textbf{NEWS En$\to$De} and \textbf{TED En$\to$It} and dropout 0.1 and label-smoothing 0.1 for \textbf{OS En$\to$De}.
This resulted in models with ca. 51M parameters for NEWS and TED and ca. 60M parameters for OS for both the sentence-level and the document-level systems.
All systems are trained until the validation perplexity does no longer improve and the best checkpoint is selected using validation perplexity as well.
Training took around 24h for NEWS and TED and around 96h for OS on a single NVIDIA GeForce RTX 2080 Ti graphics card.
Due to computational limitations, we report results only for a single run.
For the generation of segments (see Section \ref{sec:methodology}), we use beam-search on the token level with beam-size 12 and length normalization.
To calculate \BLEU \cite{papineni2002bleu} and \TER \cite{snover2006study} we use SacreBLEU \cite{post2018call}.

\end{document}